OpenStudy (anonymous):

The 3×3 matrix A is transformed into I by the following elementary row operations replace row 1 with row 1 + (4) row 3 exchange row 2 and row 3 replace row 2 with (3) row 2 exchange row 1 and row 2 exchange row 1 and row 3 a) Find the corresponding elementary matrices E1,…,E5 so that E5E4E3E2E1A=I. b) Find A.

OpenStudy (anonymous):

E1 have as \[\left[\begin{matrix}1 & 0 & 4 \\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

for E2 I got \[\left[\begin{matrix}1 & 0 & 4 \\ 0 & 0 & 1\\ 0 & 1 & 0 \end{matrix}\right]\] but they say it is wrong

OpenStudy (anonymous):

so do you know the write answer

OpenStudy (anonymous):

No I only know when I get it wrong

OpenStudy (anonymous):

okay why don't you ask the pony

OpenStudy (usukidoll):

hmmm you need an identity matrix

OpenStudy (anonymous):

i mean the UsukiDoll

OpenStudy (usukidoll):

E5E4E3E2E1A=I. what is your original matrix?

OpenStudy (usukidoll):

last two you're swapping... rows.. but err this would be easier if you could supply the original matrix

OpenStudy (usukidoll):

awww it's offline what's the point?!

ganeshie8 (ganeshie8):

sicne we have transformations performed, we can construct the original matrix A

ganeshie8 (ganeshie8):

try this for E1 : \(\left[\begin{matrix}0 & 13 & 0 \\ 1 & 0 & 0 \\ 0 & 3 & 0\end{matrix}\right]\)

ganeshie8 (ganeshie8):

try this for E2 : \(\left[\begin{matrix}0 & 1 & 0 \\ 1 & 0 & 0 \\ 0 & 3 & 0\end{matrix}\right]\)